# Forest-of-Thought: Scaling Test-Time Compute for Enhancing LLM Reasoning

### v1

@@ 3.2 Dynamic Self-Correction Strategy
To improve the probability for giving correct answer in each tree, we introduce the dynamic self-correction strategy. For the initial result $s_{i}$ from a reasoning tree, the self-correction strategy evaluates its correctness and effectiveness, and assigns corresponding scores upon completion of each reasoning step. The score can be obtained with a LLM $p_{\theta}$:
$score_{i}\sim p_{\theta}(score_{i}\mid s_{i},x).$ (7)
Once a step’s score falls below the predefined threshold, the strategy automatically triggers a correction mechanism. This mechanism first reviews and analyzes past failure cases (denoted as $\mathcal{C}$) to identify the causes of low scores and common error patterns, subsequently attempting to correct errors and optimize the reasoning direction. The correction process can be performed by the LLM itself:
$s_{i}^{\prime}\sim p_{\theta_{1}}(s_{i}^{\prime}\mid\mathcal{C},s_{i},x).$ (8)
By employing this learning-from-history and real-time correction mechanism, the model not only avoids repeating mistakes on the same issues but also finds effective solutions to new and complex problems more swiftly and accurately. This continuous learning and optimization process significantly enhances the model’s reasoning capabilities and efficiency, enabling it to perform more intelligently and effectively across a variety of complex tasks.
Especially for math tasks, we have incorporated a correction method based on mathematical rule matching defined as function $F(\cdot)$, in addition to the aforementioned mechanisms. The framework relies on predefined mathematical rules to perform rapid error correction: $s_{i}^{\prime}=F(s_{i})$, enabling it to quickly identify and rectify errors in mathematical expressions. For instance, the framework can detect common computational errors such as division by zero, mismatched parentheses, or misuse of symbols, and immediately correct them. This approach not only enhances the accuracy of the framework but also significantly reduces the time required for error detection and correction, making the entire reasoning process more efficient. The process details of the proposed dynamic self-correction strategy are shown in Algorithm 2.
@@
0:  Input context $x$, LM $p_{\theta}$, mathematical rule correction function $F$, priori knowledge sets $\mathcal{C}$;
@@
7:        $s_{i}^{\prime}\leftarrow p_{\theta}(s_{i}^{\prime}\mid\mathcal{C},s_{i},x)$;
@@ 3.3 Decision Making Strategy
To address complex mathematical problems, we designed the Consensus-Guided Expert Decision (CGED) strategy to ensure high accuracy and reliability in the final answers generated by FoT. The CGED approach combines collective intelligence with expert judgment, guiding the reasoning process from consensus-based decision-making to expert evaluation.
Selecting the Optimal Leaf Node. In the FoT method, each independent tree generates one or more possible answers through its unique reasoning path. Subtrees vote for the candidate answer that receives the most support. If a consensus cannot be reached, the math expert evaluates the reasoning processes and selects the final answer to ensure its accuracy and validity.
FoT Decision-Making Process. During reasoning, each activated tree produces an optimal solution for its path. These answers are then filtered through majority voting and expert evaluation. For complex tasks, if answers from multiple trees conflict, an LLM expert examines the reasoning processes and makes a final decision based on their expertise. This process enhances robustness by reducing errors and inconsistencies.
@@ 4.1 Experimental setups
For the Game of 24, our FoT is built using the ToT as the reasoning tree. Beyond the ToT-based FoT, we developed an MCTSr-based FoT to address more complex and diverse mathematical problems, including those in the GSM8K and MATH benchmarks.
In all the three benchmarks, the Llama3-8B-Instruct (at Meta, 2024) is used as the base language model. In addition, we extend our method on Mistral-7B (Jiang et al., 2023), and GLM-4-9B (GLM et al., 2024) on GSM8K benchmark to verify its generalization.
@@ 4.2 Game of 24
The Game of 24 originates from ToT (Yao et al., 2024), where the objective is to construct an arithmetic expression using each of the four given numbers exactly once, such that the expression evaluates to 24. The 24-point game process is shown in Figure 2. We conducted multiple inference experiments on a V100-32G GPU using Llama3-8B, optimizing parameters such as the number of trees in the forest, activation mechanisms, and sparsity settings. We evaluated the impact of different configurations on inference speed and accuracy.
Figure 2: Illustration of FoT in Game of 24 task.
@@
Results. In Table 1, we compare the results of increasing the number of leaf nodes at each layer in the ToT framework to allow for a greater diversity of potential reasoning outcomes. The hypothesis was that by expanding the leaf nodes, ToT would be able to explore more reasoning paths, leading to improved accuracy. However, when the number of leaf nodes per layer was increased to 32, the experimental results did not show significant improvement. It appears that the system reached a reasoning bottleneck, where further expansion of the leaf nodes did not lead to substantial gains in performance. This suggests that beyond a certain threshold, simply increasing the number of available reasoning paths does not necessarily translate into better reasoning outcomes, likely due to diminishing returns from excessive node expansion.
In contrast, FoT approach demonstrated a notable improvement in accuracy without a significant increase in computational cost. Specifically, when the number of trees was increased from 2 to 4, the accuracy of FoT improved by 14%, showing a significant boost in reasoning performance. Unlike the ToT method, which seemed to encounter a plateau as the number of leaf nodes increased. This suggests that the diversity of reasoning paths provided by multiple trees in FoT is more effective than simply expanding the complexity of individual trees, highlighting the advantages of the FoT framework in achieving scalable and efficient reasoning improvements.
Table 1: Performance Comparison between ToT and FoT. ToT* denotes the application of the self-correction method. Average infer times refers to the mean number of inference operations performed by the ToT or FoT across the test set, calculated as the total number of inference operations divided by the number of test instances. $b$ denotes the breadth limit of ToT, and $n$ is the number of reasoning trees.
Method Average infer times Success
IO prompt 1 3%
CoT prompt 1 3%
ToT ($b$=5) 13.74 5.26%
FoT ($b$=1, $n$=5) 14.49 9.47%
ToT* ($b$=5) 13.74 56.25%
ToT* ($b$=8) 16.82 74.71%
ToT* ($b$=16) 18.92 74.74%
ToT* ($b$=32) 22.15 76.84%
FoT* ($b$=5, $n$=2) 19.04 77.89%
FoT* ($b$=5, $n$=4) 23.64 91.58%
FoT* ($b$=5, $n$=8) 25.64 96.84%
FoT* ($b$=5, $n$=16) 26.99 100%
@@ 4.3 GSM8K Benchmark
In addition to ToT, we evaluated the benefits of integrating multiple methods into the FoT framework on the GSM8K (Cobbe et al., 2021) dataset.
Results. As shown in Figure 3, we constructed forests using various methods, including Zero-Shot-CoT, MCTSr with one turn, MCTSr with 4 rollouts, and MCTSr with 8 rollouts. The experimental results demonstrate that as the number of trees in the forest increases, the advantages of the multi-method forest approach become more pronounced. Notably, the 4-rollouts MCTSr with 2 trees achieved 3.2% higher accuracy compared to the 8-rollouts MCTSr. Additionally, it outperformed the 8-rollouts MCTSr with 2 trees, highlighting its clear advantage. These findings indicate that increasing the diversity of reasoning outcomes has a more significant impact on performance than merely extending the depth of individual trees.
When using MCTSr with 2 trees in the FoT framework, the accuracy improvement is notably significant, highlighting the substantial benefit of introducing even a small amount of diversity in reasoning paths. However, as the number of trees continues to increase, the rate of improvement gradually diminishes. This phenomenon suggests that the initial addition of trees provides a major boost by expanding the range of explored solutions, but subsequent increases contribute less incrementally to the overall accuracy. The diminishing returns can be attributed to overlapping reasoning paths or the fact that additional trees may explore regions of the solution space that are already well-covered, thereby reducing their impact. This highlights the importance of balancing tree diversity and computational efficiency to achieve optimal performance in the FoT framework.
@@
Scaling laws in FoT across various base models. In addition to the previously discussed research, we extended our exploration to evaluate the performance of different base models of similar size when reasoning with FoT framework. Specifically, we conducted experiments using three models: Mistral-7B, Llama3-8B, and GLM-4-9B. These experiments aimed to assess how the FoT scales in accuracy as the number of activated subtrees increases.
As shown in Figure 4, the results demonstrated a clear scaling law: as the number of subtrees in FoT increased, model accuracy improved significantly. This scaling behavior aligns with the theoretical expectations that activating additional subtrees enhances the diversity of reasoning paths, enabling the model to refine its problem-solving capabilities. Each additional subtree contributed incrementally to the overall performance, collectively boosting the reasoning capacity of the framework.
The observed trend suggests that the relationship between the number of activated subtrees and accuracy is both positive and predictable, adhering to scaling law principles. This consistency indicates that the FoT method efficiently utilizes computational resources to maximize reasoning accuracy. Moreover, the experiments underscore that as computational capacity is allocated to activating more subtrees, the performance gains follow a diminishing-but-consistent trajectory, demonstrating the robustness and scalability of the FoT approach.
This study not only confirms the effectiveness of FoT across various base models but also provides a theoretical foundation for optimizing resource allocation in complex reasoning tasks. It highlights the adaptability of FoT to different architectures while adhering to scaling principles, making it a versatile tool for enhancing reasoning performance in large language models.
@@ 4.4 MATH Benchmark
This section presents the results of applying the FoT algorithm across various complexity levels on the MATH dataset (Hendrycks et al., 2021). The dataset is stratified into five difficulty levels, ranging from level 1 (easiest) to level 5 (most challenging).
Results. The experimental results are shown in Table 2. The FoT (Forest of Thought) method demonstrates a consistent improvement over the MCTSr across all difficulty levels from 1 to 5 in the mathematics dataset. FoT (n=4) shows about 10% improvement from Level-1 to Level-5. This consistent enhancement highlights the effectiveness of the FoT method in handling problems of varying complexity, from the simplest to the most challenging tasks.
Table 2: FoT’s experimental findings on the math dataset. The results of our experimental reproduction are indicated by MCTSr#. The number of subtrees in FoT that are activated is denoted by n. LLaMA3-8B-Instruct is utilized in the experiments.
Data Zero-Shot MCTSr# FoT FoT
CoT ($n$=2) ($n$=4)
Level-1 57.21 67.51 73.68 77.35
437
Level-2 40.60 47.43 52.01 57.27
894
Level-3 27.32 34.22 39.35 45.09
1131
Level-4 16.64 21.66 26.28 31.05
1214
Level-5 7.10 8.38 11.10 14.50
1324
MATH 24.36 29.74 33.88 38.42
5000
@@ 4.5 Ablation Studies of Stopping Strategies
This paper compares three methods for selecting the optimal leaf node from subtrees, highlighting their effectiveness and performance:
Random Selection: A leaf node is selected randomly as the final answer. This method is simple and requires minimal computation but often results in unstable and less accurate outcomes.
Score-based Selection: The leaf node with the highest score is chosen as the final answer. The score is derived from the maximum value retained during MCTSr node expansion, reflecting the most promising path explored.
CGED Selection: This method integrates collective reasoning with expert judgment. It considers leaf node scores while incorporating expert evaluation, ensuring optimal decisions even in complex or ambiguous cases. By synthesizing insights from multiple reasoning paths, CGED enhances accuracy and reliability.
Experimental results in Table 3 show that CGED significantly outperforms the other methods, achieving 0.9% higher accuracy than Random Selection and 0.8% higher than Score-based Selection. While Random Selection offers simplicity, its randomness reduces accuracy. In contrast, CGED demonstrates greater stability and precision, particularly for complex, multi-dimensional problems.
Table 3: Subtree decision-making process. ”Random” refers to selecting a random answer from the leaf nodes as the tree’s answer, while ”Score” indicates that the answer is chosen based on the highest score among the leaf nodes. The accuracy of the experimental results on the GSM8K benchmark is displayed. The average inference times (avg infer times) typically refers to the ratio of the total number of inference steps to the number of test samples in GSM8K.
Method Accuracy Avg infer times
Random 77.73 7.00
Score 77.86 7.00
CGED 78.62 7.28
In Figure 5, we compare the forest stopping strategies based on Majority Vote, Math Expert, and CGED. In the FoT method, when two subtrees are activated, the CGED strategy demonstrates similar accuracy to the Majority Vote and Math Expert strategies. However, as the number of activated subtrees increases, the performance differences between these strategies become more pronounced. Notably, with five activated subtrees, the CGED strategy achieves a clear improvement in accuracy, outperforming the Majority Vote and Math Expert strategies by 2%. This highlights the effectiveness of the CGED method in scenarios with a higher number of activated subtrees.
@@
Figure 5: Stopping strategies in FoT. The x-axis represents the number of activated subtrees in the FoT, while the y-axis indicates the accuracy achieved on the GSM8K dataset.
4.6 Ablation Studies of Self-Correction methods
The advantages of incorporating a dynamic self-correction module into the Zero-Shot-CoT and ToT approaches are summarized in Table 4. Specifically, adding this dynamic adjustment mechanism improves the efficacy of the CoT method by 5%. Remarkably, the accuracy of the ToT approach increases from an error-prone initial state in single-step reasoning to over 50% after integrating the self-correction mechanism. This improvement is particularly significant because errors in ToT’s single-step reasoning tend to propagate through subsequent reasoning steps, severely impairing overall performance. Consequently, the dynamic self-correction process plays a critical role in ToT by promptly addressing early-stage mistakes, preventing their compounding effects, and significantly enhancing the final accuracy.
Table 4: Results of dynamic self-correction strategies. * denotes the application of the self-correction method. The experiments were conducted using the Llama-3-8B-Instruct model.
Method Benchmark Accuracy
zero-shot-cot GSM8K 74.98
zero-shot-cot* GSM8K 79.98
ToT Game of 24 5.26
ToT* Game of 24 56.25

### v2

@@ 3.2 Dynamic Self-Correction Strategy
Self-correction is a fundamental cognitive method that humans use to solve complex problems (Simon, 1991; Flower, 1981; Amabile, 1983). Unlike Self-Refine (Madaan et al., 2023), which relies on a fixed number of iterations, our approach dynamically evaluates each reasoning step, specifically by monitoring the predicted logits scores to assess the quality of the reasoning results. When the model’s score falls below a predefined threshold, a correction mechanism is automatically triggered to detect and fix errors in a timely manner. This approach enhances the flexibility and adaptability of the reasoning process, as it does not rely on preset iteration counts, but instead adjusts the model’s output based on real-time feedback.
Additionally, our method incorporates predefined mathematical rules, further improving the accuracy and reliability of the reasoning process. By embedding these rules into the reasoning framework, the model can immediately correct errors upon detection. For example, in Game 24, the model can verify whether the remaining numbers in the output are derived from the input numbers, enabling quick error detection and correction. This immediate correction mechanism not only improves reasoning quality but also significantly reduces the risk of error propagation. Overall, through dynamic evaluation, real-time feedback, and rule-driven correction mechanisms, our approach enables the model to quickly identify and correct errors, ultimately enhancing its overall performance and robustness.
The process details of the proposed dynamic self-correction strategy are shown in Algorithm 2.
@@
0:  Input context $x$, LM $p_{\theta}$, mathematical rule correction function $F$, priori knowledge sets $\mathcal{B}$;
@@
7:        $s_{i}^{\prime}\leftarrow p_{\theta}(s_{i}^{\prime}\mid\mathcal{B},s_{i},x)$;
@@ 3.3 Decision Making Strategy
To address complex mathematical problems, we designed the Consensus-Guided Expert Decision (CGED) strategy to ensure high accuracy and reliability in the final answers generated by FoT. The CGED approach combines collective intelligence with expert judgment from LLMs, guiding the reasoning process from consensus-based decision-making to expert evaluation.
FoT Decision-Making Process. During the reasoning process of FoT, each activated tree generates the optimal solution for its reasoning path. These solutions then undergo majority consensus voting and expert evaluation to gradually identify the best answer. Specifically, for complex reasoning tasks, if the majority of trees produce inconsistent results, an LLM expert will compare the reasoning processes and outcomes of the different trees, making a final decision based on their professional knowledge and experience. This approach ensures the accuracy of the final result, effectively reduces errors and biases in the reasoning process, and enhances the robustness of the entire framework.
@@ 4.1 Experimental setups
In this experimental section, we explore FoT methods based on ToT and MCTSr, conducting experiments across multiple LLMs and datasets. For the Game of 24 (Yao et al., 2024), our FoT is built using ToT as the reasoning tree. In addition to the ToT-based FoT, we developed an MCTSr-based FoT to address mathematical problems, including those from the GSM8K (Cobbe et al., 2021a) and MATH (Hendrycks et al., 2021b) benchmarks. Furthermore, we extend our method to models such as Llama3-8B-Instruct (at Meta, 2024), Mistral-7B (Jiang et al., 2023), and GLM-4-9B (GLM et al., 2024), testing it on the GSM8K benchmark to assess its generalization.
@@ 4.2 Game of 24
The Game of 24, from 4nums.com, involves constructing an arithmetic expression using each of the four given numbers exactly once, such that the expression evaluates to 24. We removed the duplicate and unsolvable problems, leaving 95 problems as the test set. In our experiment, we set the sampling temperature to 0.95. We also conducted a series of ablation experiments based on Llama3-8B-Instruct, as shown in Table 1.
Figure 2: On the Game of 24, the performance comparison between FoT and ToT, under the same computational cost.
Table 1: Ablation experiment of FoT: BoN combines the reasoning paths of two trees of thoughts, integrating their outputs to produce a final answer. In contrast, FoT activates two subtrees for reasoning.
Method Self- Input Sparse Acc. [%] LLM
Correction Enhancement Activation invoked
BoN 10.58 32.32
FoT 60.24 32.32
FoT 77.98 32.32
FoT 77.98 26.99
Results. Table 1 presents an ablation experiment comparing the performance of FoT and ToT, highlighting the impact of different components on reasoning accuracy. The experiment starts with the BoN method, which directly uses the ToT framework without input enhancement, sparse activation, or self-correction. Under this configuration, the accuracy is relatively low, at 10.58%, as the model relies solely on the base ToT approach, without any optimizations or adjustments to enhance its reasoning capabilities. In contrast, by introducing self-correction, FoT shows a significant improvement, achieving an accuracy of 60.24%. The self-correction mechanism plays a pivotal role in enhancing the reasoning process. It allows real-time adjustments to the current reasoning path, ensuring errors are corrected promptly as they are detected. This reduces the accumulation of errors that can otherwise arise when later reasoning paths depend on potentially flawed earlier steps. Consequently, FoT with self-correction delivers more accurate and reliable reasoning, especially in complex problem-solving scenarios. Further improvements are observed when input enhancement is introduced alongside self-correction, resulting in the FoT with both features. This configuration boosts the accuracy to 77.98%, highlighting the importance of enriching the model’s inputs to provide broader perspectives, which in turn supports more robust reasoning. Finally, when sparse activation is incorporated into the configuration, there is a significant reduction in the number of LLM invocations, from 32.32 to 26.99. This not only improves computational efficiency but also ensures that the model maintains its highest accuracy yet, demonstrating the power of combining self-correction, input enhancement, and sparse activation.
@@
In Figure 2, we compare the computational cost and performance of the extended ToT and FoT methods. In the experiment, we progressively increase the number of selectable leaf nodes in ToT (b = 2, 4, 8, 16, 32) to allow for greater diversity in potential reasoning outcomes. The experimental results show that simply increasing the number of leaf nodes in ToT leads to a gradual improvement in accuracy, though the gains become more modest as the number of nodes increases. However, when the number of leaf nodes per layer was increased to 32, the results showed no significant improvement. This suggests that the system reached a reasoning bottleneck, where further expansion of the leaf nodes did not result in substantial performance gains. Beyond a certain threshold, increasing the number of available reasoning paths appears to offer diminishing returns, likely due to the inefficiencies of excessive node expansion. In contrast, the FoT method, which incorporates techniques such as self-correction, input enhancement, and sparse activation, demonstrates a much more significant improvement in accuracy. These findings highlight that while expanding the structure of ToT provides only limited performance gains, the integration of additional features in FoT leads to a more substantial enhancement in model performance.
Table 2: Performance Comparison on Game of 24: Our method, with 8 activated subtrees, achieved the highest average ranking across different inference frameworks.
Method LLM invoked Success
IO 1.00 10.22%
CoT (Kojima et al., 2023) 1.00 4.38%
CoT-SC (Wang et al., 2023) 10.00 4.38%
GoT (k=1) (Besta et al., 2024a) 7.00 5.26%
ToT (b=5) (Yao et al., 2024) 13.74 74.00%
BoT (Yang et al., 2024) 3.00 82.40%
XoT (w/ 3 r) (Ding et al., 2024) 1.78 85.40%
Ours 25.64 96.84%
@@ 4.3 GSM8K Benchmark
In addition to the Game of 24, we evaluated the benefits of integrating multiple methods into the FoT framework using the GSM8K (Cobbe et al., 2021b) dataset.
Results. As shown in Figure 3, we constructed forests using various methods, including Zero-Shot-CoT, MCTSr with one turn, MCTSr with 4-rollouts, and MCTSr with 8-rollouts. The experimental results demonstrate that as the number of trees in the forest increases, the advantages of the multi-method forest approach become more pronounced. Notably, the 4-rollouts MCTSr with 2 trees achieved 3.2 % higher accuracy compared to the 8-rollouts MCTSr. Additionally, it outperformed the 8-rollouts MCTSr with 2 trees, highlighting its distinct advantage. These findings suggest that increasing the diversity of reasoning outcomes has a more significant impact on performance than simply extending the depth of individual trees.
In Table 3, we evaluate the impact of the self-correction mechanism by conducting experiments at different confidence threshold levels. We systematically adjusted the threshold and observed the model’s performance in terms of accuracy. Our findings revealed that when the threshold was set to 0.5, the model achieved significantly better accuracy compared to other threshold values. This suggests that a threshold of 0.5 strikes an optimal balance, enabling the model to identify and correct errors without being too cautious or too lenient in its self-assessment.
Table 3: The accuracy performance of the dynamic self-correction strategy with different thresholds on the GSM8K dataset.
Threshold Accuracy
0.3 87.34
0.4 88.17
0.5 90.14
0.6 88.02
@@
Scaling laws in FoT across various base models. In addition to the previously discussed research, we extended our exploration to evaluate the performance of different base models of similar size within the FoT framework. Specifically, we conducted experiments using three models: Mistral-7B, Llama3-8B, and GLM-4-9B. As shown in Figure 4, the results demonstrate a clear scaling law: as the number of activated subtrees in FoT increases, model accuracy improves significantly. This scaling behavior aligns with theoretical expectations, which suggest that activating additional subtrees enhances the diversity of reasoning paths and enables models to refine their problem-solving capabilities. Each subtree incrementally contributed to the overall performance, collectively boosting the framework’s reasoning capacity.
The observed trend suggests a predictable and positive relationship between the number of activated subtrees and accuracy, consistent with scaling law principles. This highlights that the FoT method effectively utilizes computational resources to maximize reasoning accuracy. Moreover, as computational capacity is allocated to activate more subtrees, performance gains exhibit a diminishing, yet consistent, trajectory—showcasing the robustness and scalability of FoT.
@@ 4.4 MATH Benchmark
This section presents the results of applying the FoT method across various complexity levels on the MATH (Hendrycks et al., 2021a) dataset.
Results. The experimental results are shown in Figure 5, where the performance of the FoT method is compared to MCTSr across different difficulty levels in the mathematics dataset. The results clearly demonstrate that the FoT method consistently outperforms MCTSr at every level, from Level 1 to Level 4. Specifically, FoT (n=4) exhibits an impressive and consistent improvement of approximately 10% in performance as the difficulty level increases from Level 1 to Level 4. This steady improvement highlights the robustness and versatility of the FoT method, which is able to adapt effectively to a wide range of problem complexities. At Level 1, which consists of relatively simple problems, both methods perform reasonably well. However, as the difficulty increases, FoT shows a clear advantage, achieving higher accuracy compared to MCTSr at each subsequent level. By Level 4, the most challenging problems in the dataset, the FoT method not only maintains its performance but also demonstrates a significant enhancement in problem-solving capabilities. This consistent improvement underscores the effectiveness of the FoT method in handling problems of varying complexity.
Figure 5: FoT demonstrates consistent performance across different levels of the MATH Dataset.
@@ 4.5 Ablation Studies of Stopping Strategies
In Figure 6, we present a detailed comparison of three forest stopping strategies: Majority Vote, Math Expert, and CGED. The results show the performance of these strategies across varying numbers of activated subtrees in the FoT method. When only two subtrees are activated, the accuracy of the CGED strategy is found to be quite similar to both the Majority Vote and Math Expert strategies, indicating that for simpler cases, all three strategies perform comparably. However, as the number of activated subtrees increases, the differences in performance among these strategies become more noticeable. Specifically, when five subtrees are activated, the CGED strategy demonstrates a clear improvement in accuracy, surpassing the Majority Vote and Math Expert strategies by a margin of 2%. This suggests that the CGED strategy becomes increasingly effective in handling more complex reasoning scenarios, where multiple reasoning paths are activated simultaneously. These findings underscore the superior adaptability and performance of the CGED strategy, particularly in tasks involving a larger number of activated subtrees, where it outperforms the other two strategies in terms of accuracy and overall efficiency.
@@
Figure 6: Comparison of the accuracy of FoT with three different stopping strategies across various numbers of activated subtrees: The x-axis represents the number of activated subtrees, while the y-axis shows the accuracy achieved on the GSM8K dataset for each stopping strategy.
4.6 Main Results
As shown in Table 2, the performance comparison on the Game of 24 validates the effectiveness of the FoT in the mathematical and logical reasoning task. The experimental results indicate that the FoT with 8 activated subtrees achieved the highest success rate at 96.84%, significantly outperforming other methods. In contrast, single-step reasoning methods like IO (10.22%) and CoT (4.38%) rely solely on the model’s pre-trained capabilities, showing limited performance. Multi-step reasoning techniques, such as ToT (b=5) and BoT, achieved success rates of 74.00% and 82.40%, respectively, demonstrating a notable advantage. Additionally, XoT, after three deployments, reached a success rate of 85.40%. These results clearly indicate that the FoT method, particularly with a higher number of activated subtrees, enhances the model’s self-correction capabilities and significantly improves the performance on the Game of 24 while only modestly increasing the computational cost.
The results presented in Table 4 highlight the performance of the FoT method across various datasets, including GSM8K, MATH-500 (Lightman et al., 2023), and AIME 2024 (AI-MO, 2024). In the GSM8K dataset, FoT with 4 activated subtrees (FoT ($n$=4)) based on QwQ achieved the accuracy of 97.33%, outperforming rStar-Math (95.20%) and GPT-4o (92.90%). FoT ($n$=4) with QwQ achieved a strong performance with an accuracy of 91.20%, surpassing rStar-Math64 (90.00%). On AIME 2024, FoT showed significant improvements, reaching 53.33%, which is much higher than rStar-Math at 6.66%. These results underscore the effectiveness of FoT in enhancing model performance, particularly as the number of activated subtrees increases, demonstrating its potential to improve reasoning accuracy across a range of mathematical problem-solving tasks.
Table 4: The following summarizes the performance of FoT and other state-of-the-art language models across different levels of mathematical benchmark tests. GSM-level refers to the GSM8K, MATH-level to the MATH500, and Olympiad-level to AIME 2024. Specifically, rStar-Math (Guan et al., 2025) refers to rStar-Math (7B SLM + 7B PPM), and the latter denotes the Pass@1 accuracy achieved when sampling 64 trajectories.
Model GSM-level MATH-level Olympiad-level
GPT-4o 92.90 76.60 9.30
rStar-Math 95.00 89.40 46.70
rStar-Math64 95.20 90.00 53.30
Base Model: Qwen2.5-Math-7B-Instruct
Qwen2.5-7B 88.48 82.60 6.00
FoT ($n$=2) 93.33 83.02 26.67
FoT ($n$=4) 95.00 85.80 33.33
FoT ($n$=8) 96.89 86.20 46.67
Base Model: QwQ-32B-preview
QwQ 95.30 90.60 50.00
FoT ($n$=4) 97.33 91.20 53.33